OpenStudy (anonymous):

solve ; 113000/(x+235) when y=200

OpenStudy (anonymous):

\[200=\frac{113.000}{x+235}\]?

OpenStudy (anonymous):

200*(x+235)=113.000\[x+235=\frac{113.000}{200}\]\[x=565-235\] x=330